The ratios of distances from point p (x,y) to A (4,5)  and B(5,4) is 2:3 then find equations of locus
 

Grade:10

1 Answers

Arun
5 years ago
Dear Nikhil
 
PA/PB = 2/3
 
\sqrt(x-4)² + (y-5)² / \sqrt(x -5)² + (y -4)²  = 2/3
9 [(x-4)² + (y-5)²] = 4 [(x -5)² + (y-4)²]
 
Open it. 
This will be locus of P